In a synthesis experiment 3mL of 25% (v/v) en in ethanol was used to make a complex Nickel salt. Calculate the mols of en used in this synthesis. You may assume a density of ethanol of 0.789g/mL and a density for en of 0.899g/mL. Report your answer in units of mol.
Answer #1

Since we have a 25% v/v solution of en in ethanol, and we use 3 ml of solution; the volume of pure en in this solution would be:

25/100 * 3 = 0.75 ml

The weight of this 0.75 ml of en would be: volume * density = 0.75 * 0.899 g = 0.674 g

Moles of en = weight / molar mass = 0.674 / 60.1 = 0.0112 moles